# Odd-parity Magnetism Driven by Antiferromagnetic Exchange

### v2

@@ Abstract
Induced odd-parity magnetism: Here, two key ingredients are required to induce odd-parity spin splittings. First, the AFM state must belong to an irreducible representation (IR) that is at least two-dimensional (2D) so that there are two magnetic degrees of freedom. Second, the inversion operator for this IR must be non-trivial. Both ingredients are natural consequences of non-symmorphic space groups.
These two symmetry ingredients are illustrated in the simplest unit cell doubling AFM in Fig. 1. The nonmagnetic unit cell consists of two sublattice atoms. The two sublattice magnetic moments, here labeled ${\bf S_{1}}$ and ${\bf S_{2}}$, typically form a 2D IR for AFM in non-symmorphic space groups. As described below, consideration of the inversion operator naturally leads to two classes. In the class 1 (top panel), the two atoms are related by inversion, whereas in the class 2 (bottom panel), each atom sits at an inversion center.
Here ${\bf S_{1}}\times{\bf S_{2}}$ is an important quantity, as we discuss later, it generates the odd-parity spin-splitting. This induced order is time-reversal invariant, as time-reversal symmetry flips both spins, and is translationally invariant. The inversion symmetry acts non-trivially on ${\bf S_{1}}$ and ${\bf S_{2}}$ (into blue dashed arrows). In the class 1, inversion interchanges ${\bf S_{1}}$ and ${\bf S_{2}}$. In the class 2, inversion leaves ${\bf S_{1}}$ unchanged but flips ${\bf S_{2}}$. In both cases, the resulting spin splitting ${\bf S_{1}}\times{\bf S_{2}}$ is odd under inversion. Here ${\bf S_{1}}\times{\bf S_{2}}$ can also be interpreted as a non-relativistic version of the Dzyaloshinsky-Moriya vector Dzyaloshinsky (1958); Moriya (1960).
@@
It is possible to identify all space groups and AFM ordering wavevectors that give rise to the two required symmetry ingredients. The nontrivial inversion operator requires non-symmorphic space groups, where inversion does not commute with all symmetries at certain time-reversal and inversion invariant momenta (TRIM). This includes 325 IRs for $p$-wave, 84 IRs for $f$-wave, and 12 IRs for $h$-wave spin splittings. The list of space groups and wavevectors is in SM. In these 421 distinct scenarios, the AFM state with odd-parity magnetism can be stabilized through a single continuous phase transition. These 2D IRs are not restricted to cases with only two atoms per nonmagnetic unit cell. For systems with more atoms, the theory remains applicable. However, $\bf S_{1}$ (and $\bf S_{2}$) now describes spin order across multiple atoms in the non-magnetic unit cell. The $p$-, $f$-, and $h$-wave spin splittings are identified by analyzing other symmetry operators at these TRIMs.
Figure 2: (Left) FeSe coplanar magnetic state (Center) DFT band dispersion along the momentum cut ±$\pm(\pi,0.4\pi,0.5\pi)$. (Right) Fermi surface ($k_{z}=0.4\pi$ section) exhibiting $h$-wave spin splitting. The band spin induced is represented by a red-white-blue color code.
To test our symmetry-based arguments, we have examined if the predicted $h$-wave $k_{x}k_{y}k_{z}(k_{x}^{2}-k_{y}^{2})\sigma_{z}$ spin-splitting appears within a DFT calculation for FeSe with coplanar order at the M-point $(\pi,\pi,0)$. Collinear AFM ordering at the M point is known to generate the nematic order observed in many Fe-based materials Kreisel et al. (2020). A coplanar magnetic order is sometimes observed in this class of materials Ding et al. (2017); Stadel et al. (2022). In particular, this coplanar state appears in LaFeAs1-xPx Stadel et al. (2022) with space group 129, providing a materials realization of this state. As shown in Fig. 2, the DFT results for this coplanar state agree with our symmetry prediction. Furthermore, this calculation reveals that the spin-splitting, even for a highly nodal $h$-wave state, can be as large as 0.1 eV.
@@
Here, we have included all the symmetry-allowed terms to quartic order in $\bf S_{1,2}$. The continuous phase transition occurs as $\alpha(T)\rightarrow 0$. Depending on $\beta_{i}$, there are three competing AFM states with different uniform induced orders (Fig. 3).
@@
$H=\begin{pmatrix}\epsilon_{k}+t\cos\frac{k_{x}}{2}\tau_{x}&O_{M}\\ O_{M}^{\dagger}&\epsilon_{k+Q}-t\sin\frac{k_{x}}{2}\tau_{y}\end{pmatrix}.$ (2)
Here $\tau_{i}$ are Pauli matrices for nonmagnetic sublattices and the on-site exchange driven spin ordering is described by $O_{M}=\vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}$. As in the usual SDW theoryAuerbach (2012), $H$ features sublattice-independent hopping $\epsilon_{k}$ and $\epsilon_{k+Q}$ on the diagonal blocks. The model also exhibits inter-sublattice hopping $t_{k}\tau_{x}$ and $t_{k+q}\tau_{y}$ on the diagonal ($t_{k}=t\cos\frac{k_{x}}{2}$). The transformation from $\tau_{x}$ to $\tau_{y}$ is a consequence of our choice of Fourier basis. In the nonmagnetic unit cells, sites $a_{2}$ and $b_{2}$ are equivalent, and their hopping to site $a_{1}$ can give the real term $t(e^{ik_{x}/2}+e^{-ik_{x}/2})\tau_{x}=t\cos\frac{k_{x}}{2}\tau_{x}$. In the doubled unit cell, since $a_{2}$ and $b_{2}$ are inequivalent, these two hoppings are described by different operators, resulting in both real ($t\cos\frac{k_{x}}{2}\tau_{x}$) and imaginary ($t\sin\frac{k_{x}}{2}\tau_{y}$) terms. The transformation from $\tau_{x}$ to $\tau_{y}$ can also be understood as a consequence of time-reversal symmetry that exists in the nonmagnetic state: the odd time-reversal symmetry of $\sin\frac{k_{x}}{2}$ can only be compensated by transforming $\tau_{x}$ to $\tau_{y}$.
@@
where $t_{1}\equiv(t_{x}+\widetilde{t_{x}})/2$, $t_{2}\equiv(t_{y}+\widetilde{t_{y}})/2$, $t_{3}\equiv(t_{x}-\widetilde{t_{x}})/2$, and $t_{4}\equiv(t_{y}-\widetilde{t_{y}})/2$. The Pauli matrices $\tau_{i}$ and $\rho_{i}$ act in sublattices $(1,2)$ and $(a,b)$, respectively. The time-reversal operator and inversion operator are $T=i\sigma_{y}K$ and $I=\tau_{x}$, with momentum flip. The first block, $h_{0}({\bf k})$ is the same as the nonmagnetic Hamiltonian, containing k-even $\tau_{x}$ and k-odd $\tau_{y}$ hopping terms, whose detailed form differs in different nonsymmorphic space groups (listed in SM). $\widetilde{h_{0}}({\bf k+Q})$ is obtained from $h_{0}({\bf k+Q})$ by correcting $\tau_{x,y}$ under time-reversal symmetry. The $\epsilon_{0}$ (or $t_{0}$) term includes the even (or odd) intra-sublattice hopping terms under $\bf k\rightarrow k+Q$, such as $\cos k_{y}$ (or $\cos k_{x}$) in the pedagogical example. As a next-nearest neighbor hopping, $t_{0}$ term does not significantly affect the magnitude of the odd-parity spin splitting, and the subsequent analysis assumes $t_{0}=0$. The three AFM states are described by $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},\hat{x}),J(\hat{x},0),\text{and }J(\hat{x},\hat{y})$. We note that this Hamiltonian can always be block-diagonalized since $[\rho_{z}\sigma_{z},H]=0$, allowing for analytical results.
@@
We now consider SG129(P4/nmm) with Wyckoff position 2a, relevant for FeSe (and 2c, for CeNiAsO Birk Hellenes et al. (2023)). The coefficients in the Hamiltonian are: $t_{x}=t_{x0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$ and $t_{y}=t_{y0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$ (for 2c, remove $\cos k_{x}-\cos k_{y}$ factors here and below). For the AFM state with ${\bf Q}=(\pi,0,0)$, the other coefficients are $\widetilde{t_{x}}=-t_{y0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$ and $\widetilde{t_{y}}=-t_{x0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$, from which we get $p$-wave spin splitting proportional to $\sin k_{x}$. For the AFM state at $(\pi,\pi,0)$, the other coefficients are $\widetilde{t_{x}}=t_{x0}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}$ and $\widetilde{t_{y}}=-t_{y0}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$, from which we obtain an $h$-wave spin splitting proportional to $\sin k_{x}\sin k_{y}(\cos k_{x}-\cos k_{y})\sin k_{z}$ found in Fig. 2.
@@
Spin-scalar odd-parity state: We now consider the scalar odd-parity state. We will focus on state 1 in class 1 below, and the analysis for state 2 in class 2 is provided in SM. Although the state $(\vec{J}_{1},\vec{J}_{2})=J(\hat{x},\hat{x})$ has no spin-splitting [see Eq. (4)] and does not exhibit any symmetry breaking in the electronic dispersion, it can generate a non-relativistic Berry curvature dipole (unlike the spin-vector or nematic states), strongly contributing to nonlinear I-V transport. The Berry curvature is:
@@
The derivation is in the SM. $\epsilon_{abc}$ is the antisymmetric tensor. The nontrivial k-odd dependence from $t_{x}t_{y}$ carries the site symmetry, as the AFM state distinguishes the two sublattices (state 1 in Fig. 3). This k-odd Berry curvature can generate non-zero Berry curvature dipoles $\langle d\Omega^{ij}/dk_{m}\rangle_{FS}$, which drives non-relativistic nonlinear I-V transport coefficients Sodemann and Fu (2015); Kaplan et al. (2024). In addition, this state is related to multiferroicity Cheong and Mostovoy (2007); Matsuda et al. (2024) and will also exhibit ferroelectric-like orderCheong and Mostovoy (2007).
@@ Appendix C Class 1
The three AFM states are described by $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},\hat{x}),J(\hat{x},0),\text{and }J(\hat{x},\hat{y})$. Without loss of generality, we will consider $\vec{J}_{1,2}$ in the xy plane. Using $[\rho_{z}\sigma_{z},H]=0$, we can block-diagonalize the Hamiltonian:
$\begin{split}&u^{\dagger}Hu=\left(\begin{array}[]{cc}H_{\alpha=+}&\\ &H_{\alpha=-}\end{array}\right),\;u=\left(\begin{array}[]{cccc}1&&&\\ &&1&\\ &&&1\\ &1&&\end{array}\right),\\ &H_{\alpha}=t_{1}\widetilde{\rho_{0}}\tau_{x}+t_{2}\widetilde{\rho_{0}}\tau_{y}+t_{3}\widetilde{\rho_{z}}\tau_{x}+t_{4}\widetilde{\rho_{z}}\tau_{y}+\vec{J}_{1\alpha}\cdot\vec{\widetilde{\rho}}+\tau_{z}\vec{J}_{2\alpha}\cdot\vec{\widetilde{\rho}},\end{split}$ (7)
@@
The eigenvectors $[v_{1},v_{2}]^{T}$ of $\widetilde{H}_{\alpha}$ satisfy:
@@
$\begin{split}M^{\dagger}M&=|\vec{J}_{1}|^{2}+|\vec{J}_{2}|^{2}+t_{1}^{2}+t_{2}^{2}+t_{3}^{2}+t_{4}^{2}+2(t_{1}J_{1x}+\alpha t_{2}J_{1y}+t_{3}J_{2x}+\alpha t_{4}J_{2y})\widetilde{\tau_{1}}\\ &+2(\alpha t_{1}J_{1y}-t_{2}J_{1x}+\alpha t_{3}J_{2y}-t_{4}J_{2x})\widetilde{\tau_{2}}+2(-\vec{J}_{1}\cdot\vec{J}_{2}+t_{1}t_{3}+t_{2}t_{4})\widetilde{\tau_{3}}\\ MM^{\dagger}&=|\vec{J}_{1}|^{2}+|\vec{J}_{2}|^{2}+t_{1}^{2}+t_{2}^{2}+t_{3}^{2}+t_{4}^{2}+2(t_{1}J_{1x}-\alpha t_{2}J_{1y}-t_{3}J_{2x}+\alpha t_{4}J_{2y})\widetilde{\tau_{1}}\\ &+2(\alpha t_{1}J_{1y}+t_{2}J_{1x}-\alpha t_{3}J_{2y}-t_{4}J_{2x})\widetilde{\tau_{2}}+2(+\vec{J}_{1}\cdot\vec{J}_{2}+t_{1}t_{3}+t_{2}t_{4})\widetilde{\tau_{3}}.\end{split}$ (10)
@@
The contribution from $v_{1}$ (or $v_{2}$) can be determined from $MM^{\dagger}$ (or $M^{\dagger}M$). For a general $2\times 2$ Hamiltonian $H=h_{1}\tau_{x}+h_{2}\tau_{y}+h_{3}\tau_{z}$, the Berry curvature is $\Omega_{ij}=\frac{1}{2h^{3}}\sum_{a,b,c=1}^{3}\epsilon_{abc}h_{a}\partial_{i}h_{b}\partial_{j}h_{c}$. We can then apply this formula to obtain the contribution from $v_{1}$ and $v_{2}$, with an extra prefactor of $1/4$ from their normalization. For the nematic state $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},0)$ and odd-parity spin-splitting state $(\vec{J}_{1},\vec{J}_{2})=J(\hat{x},\hat{y})$, these two contribution exactly cancel each other, leading to zero Berry curvature. For the scalar odd-parity state $(\vec{J}_{1},\vec{J}_{2})=J(\hat{x},\hat{x})$, the Berry curvature is
@@ Appendix D Class 2
The three AFM states are described by $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},\hat{x}),J(\hat{x},0),\text{and }J(\hat{x},\hat{y})$.Without loss of generality, we will consider $\vec{J}_{1,2}$ in the xy plane. Using $[\rho_{z}\sigma_{z},H]=0$, we can block-diagonalize the Hamiltonian:
$\begin{split}&u^{\dagger}Hu=\left(\begin{array}[]{cc}H_{\alpha=+}&\\ &H_{\alpha=-}\end{array}\right),\;u=\left(\begin{array}[]{cccc}1&&&\\ &&1&\\ &&&1\\ &1&&\end{array}\right),\\ &H_{\alpha}=t_{1}\widetilde{\rho_{0}}\tau_{x}+t_{2}\widetilde{\rho_{0}}\tau_{y}+t_{3}\widetilde{\rho_{z}}\tau_{x}+t_{4}\widetilde{\rho_{z}}\tau_{y}+t_{5}\widetilde{\rho_{0}}\tau_{z}+t_{6}\widetilde{\rho_{z}}\tau_{z}+t_{0}\widetilde{\rho_{z}}+\vec{J}_{1\alpha}\cdot\vec{\widetilde{\rho}}+\tau_{z}\vec{J}_{2\alpha}\cdot\vec{\widetilde{\rho}},\end{split}$ (16)

### v3

@@ Abstract
Induced odd-parity magnetism: Here, two key ingredients are required to induce odd-parity spin splittings. First, the AFM state must belong to an irreducible representation (IR) that is at least two-dimensional (2D) so that there are two magnetic degrees of freedom. Second, the inversion operator for this IR must be non-trivial. As discussed in more detail below, both ingredients are natural consequences of non-symmorphic space groups.
These two symmetry ingredients are illustrated in the simplest unit cell doubling AFM in Fig. 1. The non- magnetic unit cell consists of two sublattice atoms. The two sublattice magnetic moments, here labeled $\bf S_{1}$} and $\bf S_{2}$}, typically form a 2D IR for AFM Elcoro et al. (2017). As described below, consideration of the inversion operator naturally leads to two classes. In the class 1 (top panel), the two atoms are related by inversion, whereas in the class 2 (bottom panel), each atom sits at an inversion center.
Here ${\bf S_{1}}\times{\bf S_{2}}$ is an important quantity, as we discuss later, it generates the odd-parity spin-splitting. This induced order is time-reversal invariant, as time-reversal symmetry flips both spins, and is translationally invariant. Here, the inversion symmetry in non-symmorphic space groups acts non-trivially in the 2D IR Elcoro et al. (2017). In class 1, inversion interchanges $\bf S_{1}$} and $\bf S_{2}$}. In class 2, inversion leaves $\bf S_{1}$} unchanged but flips $\bf S_{2}$}. In both cases, the resulting spin splitting ${\bf S_{1}}\times{\bf S_{2}}$ is odd under inversion. Here ${\bf S_{1}}\times{\bf S_{2}}$ can also be interpreted as a non-relativistic version of the Dzyaloshinsky-Moriya vector Dzyaloshinsky (1958); Moriya (1960). In contrast, inversion symmetry within symmorphic space groups is always a trivial identity operation for any IR because it commutes with all other symmetry operations. Realizing odd-parity ${\bf S_{1}}\times{\bf S_{2}}$ then requires multiple IRs/phase transitions (See examples in End Matter).
@@
It is possible to identify all space groups and AFM ordering wavevectors that give rise to the two required symmetry ingredients. The nontrivial inversion operator requires non-symmorphic space groups, where inversion does not commute with all symmetries at certain time-reversal and inversion invariant momenta (TRIM) Szabó and Ramires (2024). This includes 325 IRs for $p$-wave, 84 IRs for $f$-wave, and 12 IRs for $h$-wave spin splittings. The list of space groups and wavevectors is in supplementary materials sup (including Ref. Weinert et al. (2009); Koelling and Harmon (1977); MacDonald et al. (1980); Perdew et al. (1996); Villars (1997); Shishidou et al. (2018); Koepernik and Eschrig (1999)). In these 421 distinct scenarios, the AFM state with odd-parity magnetism can be stabilized through a single continuous phase transition. These 2D IRs are not restricted to cases with only two atoms per nonmagnetic unit cell. For systems with more atoms, the theory remains applicable. However, $\bf S_{1}$ (and $\bf S_{2}$) now describes spin order across multiple atoms in the non-magnetic unit cell. The $p$-, $f$-, and $h$-wave spin splittings are identified by analyzing other symmetry operators at these TRIMs.
Figure 2: (Left) FeSe coplanar magnetic state (Center) DFT band dispersion along the momentum cut $\pm(\pi,0.4\pi,0.5\pi)$. (Right) Fermi surface ($k_{z}=0.4\pi$ section) exhibiting $h$-wave spin splitting. The band spin induced is represented by a red-white-blue color code.
To test our symmetry-based arguments, we have examined if the predicted $h$-wave $k_{x}k_{y}k_{z}(k_{x}^{2}-k_{y}^{2})\sigma_{z}$ spin-splitting appears within a DFT calculation for FeSe with coplanar order at the M-point $(\pi,\pi,0)$. Collinear AFM ordering at the M point is known to generate the nematic order observed in many Fe-based materials Kreisel et al. (2020). A coplanar magnetic order is sometimes observed in this class of materials Ding et al. (2017); Stadel et al. (2022). In particular, this coplanar state appears in LaOFeAs1-xPx Stadel et al. (2022) with space group 129, providing a materials realization of this state. As shown in Fig. 2, the DFT results for this coplanar state agree with our symmetry prediction. Furthermore, this calculation reveals that the spin-splitting, even for a highly nodal $h$-wave state, can be as large as 0.1 eV.
@@
Here, we have included all the symmetry-allowed terms to quartic order in $\bf S_{1,2}$. In $f$, spin rotational invariance allow terms ${\bf S}_{1}\cdot{\bf S}_{2}$ and $|{\bf S}_{1}|^{2}-|{\bf S}_{2}|^{2}$ to also appear, here non-symmorphic symmetries are important since they enforce that these terms are not permitted. The continuous phase transition occurs as $\alpha(T)\rightarrow 0$. Depending on $\beta_{i}$, there are three competing AFM states with different uniform induced orders (Fig. 3).
@@
$\begin{split}H&=\begin{pmatrix}h_{0}({\bf k})&O_{M}\\ O_{M}^{\dagger}&\widetilde{h_{0}}({\bf k+Q})\end{pmatrix}\\ &=\begin{pmatrix}\epsilon_{k}+t\cos\frac{k_{x}}{2}\tau_{x}&O_{M}\\ O_{M}^{\dagger}&\epsilon_{k+Q}+t\sin\frac{k_{x}}{2}\tau_{y}\end{pmatrix}\end{split}$ (2)
Here $\tau_{i}$ are Pauli matrices for nonmagnetic sublattices. The on-site exchange driven spin ordering is described by $O_{M}=\vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}$. As in the usual SDW theoryAuerbach (2012), $H$ features sublattice-independent hopping $\epsilon_{k}$ and $\epsilon_{k+Q}$ on the diagonal blocks. The model also exhibits inter-sublattice hopping $\tau_{x,y}$. The differences between $h_{0}({\bf k})$ and $\widetilde{h_{0}}({\bf k+Q})$ in Eq. (2) arise from the Q-originated Bloch phase: In $h_{0}({\bf k})$, the hopping from $(a_{2},b_{2})$ to $a_{1}$ is real, $t/2(e^{ik_{x}/2}+e^{-ik_{x}/2})=t\cos\frac{k_{x}}{2}$. In $\widetilde{h_{0}}({\bf k+Q})$, however, the relative phase factor of 1-1 between the $a$ and $b$ cells arising from $\mathbf{Q}$ causes this hopping to be imaginary, $t/2(-e^{ik_{x}/2}+e^{-ik_{x}/2})=-it\sin\frac{k_{x}}{2}$, and then to account for the resulting factor of $i$, $\tau_{x}$ is transformed into $\tau_{y}$. The transformation from $\tau_{x}$ to $\tau_{y}$ can also be understood as a consequence of time-reversal symmetry that exists in the nonmagnetic state: the odd time-reversal symmetry of $\sin\frac{k_{x}}{2}$ can only be compensated by transforming $\tau_{x}$ to $\tau_{y}$.
@@
[…]
@@
We now consider SG129(P4/nmm) with Wyckoff position 2a, relevant for FeSe (and 2c, for CeNiAsO Birk Hellenes et al. (2023)). The coefficients in the Hamiltonian are: $t_{x}=t_{x0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$ and $t_{y}=t_{y0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$ (for 2c, remove $\cos k_{x}-\cos k_{y}$ factors here and below). For the AFM state with ${\bf Q}=(\pi,0,0)$, the other coefficients are $\widetilde{t_{x}}=-t_{y0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}({-}\cos k_{x}-\cos k_{y})\sin k_{z}$ and $\widetilde{t_{y}}=t_{x0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$, from which we get $p$-wave spin splitting proportional to $\sin k_{x}$. For the AFM state at $(\pi,\pi,0)$, the other coefficients are $\widetilde{t_{x}}=t_{x0}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}$ and $\widetilde{t_{y}}=-t_{y0}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$, from which we obtain an $h$-wave spin splitting proportional to $\sin k_{x}\sin k_{y}(\cos k_{x}-\cos k_{y})\sin k_{z}$ found in Fig. 2.
@@
[…]
@@
The derivation is in the SM. $\epsilon_{abc}$ is the antisymmetric tensor. The nontrivial k-odd dependence from $t_{x}t_{y}$ carries the site symmetry, as the AFM state distinguishes the two sublattices (state 1 in Fig. 3). This k-odd Berry curvature can generate non-zero Berry curvature dipoles $\langle d\Omega^{ij}/dk_{m}\rangle_{FS}$, which drives non-relativistic nonlinear I-V transport coefficients Sodemann and Fu (2015); Kaplan et al. (2024); Watanabe et al. (2024). In addition, this state is related to multiferroicity Cheong and Mostovoy (2007); Matsuda et al. (2024) and will also exhibit ferroelectric-like orderCheong and Mostovoy (2007).
@@ Appendix C Class 1
[…]
@@
The eigenvectors $v_{1},v_{2}]^{T}$ of $\widetilde{H}_{\alpha}$ satisfy:
@@
[…]
@@
[…]
@@ Appendix D Class 2
[…]